CRY+VIS Interface Meeting
Date: 2022/7/8 13:00-13:30
Participants: Takahashi, Uchiyama, Yokozawa, Ushiba, Kimura, Yasui, Aso, Hirata, Akutsu, Ikeda, Sato, Miyoki, Node
IX: Offload of the F0. Adjustment of the OpLev for MN/PF.
EX: Adjustment of the OpLev for MN/PF.
- EY: Leak test (Leak point was in the GV itself). Repair of the GV. Offload of the GAS filters.
- Type-B: Replacement of the mirror in SRM.
Type-Bp: Adjustment of the OpLev in PR2/PR3. Investigation of the PR2 suspension. Health check in the next week.
- Type-C: Health check of the IMMT1/2 in the next week.
- Replacement of cables: EXT and BF temp.
- TM height setting:
- Commissioning in air:
- Commissioning in vacuum: Keeping the X-arm. Health check in Y-arm.
- Acceptance check:
- New vacuum pumps: When should we implement the ion pumps? It is better to do before cooling.
- Heater control: Pending.
Date: 2022/7/8 14:00-14:15
Participants: Takahashi, Uchiyama, Fabian, Tanaka, Aso, Ushiba, Washimi
- Offloaded the GAS filters after the pump down in EY.
- Replaced the cables for BF-FR, BF-TEMP, BF-GAS in EY.
- We replaced the small SRM mirror. The process went smoothly. I measured IM transfer functions yesterday, but I haven't compared with older measurements. Klog pending.
- With Yokozawa-san, we checked whether the SRM and SR3 oplev beams were reaching the central regions of their respective mirrors. The beam in SR3 is a little lower than expected, but not by too much.
- Visual inspection of PRs. PR2 only has a small hexagon between the BF and the IM.
- Type-B: Health check.
- Type-Bp: Health check.
- Type-C: Health check of IMMT1/2.
(Date in Kamioka)
The week from 7/11
- Hirata: 7/12-15
The week from 7/18